OpenStudy (anonymous):

f(x) = cos ^x theta + sin ^x theta , find 3f(4)- 2f(6)

OpenStudy (anonymous):

\[f(x)=\cos^x\theta+\sin^x\theta\] For \(f(4)\), wherever you see an \(x\) in the above function, you replace it with a 4: \[f(4)=\cos^4\theta+\sin^4\theta\] \(f(6)\) is found similarly; replace \(x\) with 6.

OpenStudy (anonymous):

and then? i understood that part, but i don't know what to do next